OpenStudy (anonymous):

help!! 1 medal and fan In sunlight, a vertical stick 6 ft tall casts a shadow 2 ft long. At the same time a nearby tree casts a shadow 14 ft long. How tall is the tree? Round to the nearest tenth. a.42.0 ft b.18.0 ft c. 85.7 ft d. 52.5 ft

OpenStudy (anonymous):

oops lets do it correctly \[\frac{6}{2}=\frac{x}{14}\]

OpenStudy (anonymous):

or \(3=\frac{x}{14}\) making \(x=14\times 3\)

OpenStudy (anonymous):

x=42

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

thanks